you have 6 subjects how do you randomly assign 3 subjects to group 1 and 3 subjects to group 2?
Answer by jim_thompson5910(35256)

There are 6 C 3 = (6!)/(3!(6-3)!) = (720)/(6*6) = 20 different ways to assign 3 subjects into group 1. Since there are 3 subjects left, they automatically go to group 2


So this means that there are 20 different ways to do this.

Note: this is assuming that order doesn't matter